a car went 346 miles in 6 hours. part of the time the car traveled at 45 mph and part of the time the car traveled at 64 mph. How long did the car travel at each speed?
Found 2 solutions by josgarithmetic, amarjeeth123:
Answer by josgarithmetic(39617)

RT=D relates rate time distance/


____________________speed_____________time___________distance
____________________45_________________u_____________45u
____________________64_________________v_____________64v
Totals_________________________________6_____________346

Answer by amarjeeth123(569)

Let the time the car traveled at 45 mph be x.
Then the time the car traveled at 64 mph is (6-x).
Distance traveled at 45 mph=45x
Distance traveled at 64 mph=64(6-x)
Total distance covered=346 miles
Distance traveled at 45 mph+Distance traveled at 64 mph=Total distance covered
45x+64(6-x)=346
45x+384-64x=346
64x-45x=384-346
19x=38
x=2 hours
The car traveled for 2 hours at 45 mph and 4 hours at 64 mph.